OpenStudy (anonymous):

FIND THE VERTEX please help f(x)=x^2-10-3

OpenStudy (anonymous):

I know that a =1 and b=-10 so then it would look like this \[x=\frac{ -10 }{ 2*(1) } =5\]

OpenStudy (anonymous):

y=f(5) y=(5)^2-10*5-3 then i will simplify y=5^2-10*5-3 y=25-10*5-3 RIGHT? SO THEN 25-10=15 AND 5-3 IS 2 RIGHT SO THEN THE GIVEN PARABOLA IS (5, ??/)

OpenStudy (anonymous):

y=f(5) y=(5)^2-10*5-3 then i will simplify y=5^2-10*5-3 y=25-10*5-3 RIGHT? SO THEN 25-10=15 AND 5-3 IS 2 RIGHT SO THEN THE GIVEN PARABOLA IS (5, ??/)